OpenStudy (anonymous):

A tourist being chased by an angry bear is running in a straight line toward his car at a speed of 4.2 m/s. The car is a distance d away. The bear is 30 m behind the tourist and running at 6.0 m/s. The tourist reaches the car safely. What is the maximum possible value for d?

OpenStudy (anonymous):

the maximum possible value for d occurs when tourist and the bear reach the car at the same time . now distance=velocity *time let the tourist cover the distance in a time t then d= 4.2*t now the bear also takes the same time to reach the car the d+30=6*t d/4.2 = (d+30)/6 then solve for d you will get d = 70m